OpenStudy (anonymous):

In a geometric progression, the second term has 9 less than the first term. The sum of the second and third terms is 30. Given that all the terms of the progression are positive, find the first term. -Trying to solve this problem for awhile now. Also, am confused when to use Sn = a(r^n-1)/r-1 or Sn = a(1-r^n)/1-r. Any help would be appreciated, thanks.

OpenStudy (anonymous):

first term is a=x second term, ar=x-9 third term, ar^2=30 r=x-9/x a*r^2 shall give u the answer

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (anonymous):

@Saarthakjain Hey, thanks for the reply. I don't understand how the third term is: ar^2=30? The question mentions the sum of the second and third terms = 30. Does that mean that S3=30-x? (x being the first term). @YanaSidlinskiy Thanks!

OpenStudy (yanasidlinskiy):

Your Welcome:D

OpenStudy (amistre64):

\[\frac{-1}{-1}*\frac{1-r^n}{1-r}=\frac{r^n-1}{r-1}\] so your Sn formulas are the same thing.

OpenStudy (amistre64):

the sum of the second and third terms is not properly a partial sum. partial sums are when we add all the terms from 1 to N

OpenStudy (amistre64):

a = a ar = a-9 ar^2 = 39 - a seems to be what we have to work with

OpenStudy (anonymous):

Yup, I got that far. Not sure how to go from there though.

OpenStudy (amistre64):

justa few thoughts 9 < a < 39 0 < a-9 < 30 and r < 1, since ar < a

OpenStudy (amistre64):

r = 1 - 9/a ar = sqrt(a*ar^2) = sqrt(39a - a^2) r = sqrt(39a - a^2)/a since r=r (a-9)/a = sqrt(39a - a^2)/a a-9 = sqrt(39a - a^2) (a-9)^2 = 39a - a^2 a^2 -18a +81 = 39a - a^2 2a^2 -57a +81 = 0

OpenStudy (anonymous):

Ok. I think I got it: Assuming that the ratio is constant, \[(a-9)/(a) = (39-1)/(a-9)\] \[a^2-18a+81 = 39a - a^2\] \[2a^2-57a+81 = 0\] \[\therefore a = 81\]

OpenStudy (amistre64):

seems like we concur :)

OpenStudy (anonymous):

Awesome, thanks for the enlightenment!! :)

OpenStudy (amistre64):

youre welcome, and good job as well :)

OpenStudy (amistre64):

i get a=27 a = 27 ar = 18 ar^2 = 12

OpenStudy (anonymous):

Oops, sorry. I got 21 as well. Somehow I typed 81 instead.

OpenStudy (amistre64):

i typo all the time, its these fat stubby fingers on this tiny little keyboard fer sure :)

OpenStudy (anonymous):

Haha, is there anyway to edit something I've posted?

OpenStudy (amistre64):

no, theres not :) you simply have to clarify it

OpenStudy (amistre64):

you asked about trig in the message, any specific example you have to work with?